A random variable is normally distributed with a mean of 20 and a standard deviation of 4. If an observation is randomly selected from the distribution, a. What value will be exceeded 15% of the time? b. What value will be exceeded 75% of the time? c. Determine two values of which the smaller has 20% of the values below it and the larger has 20% of the values above it. d. What value will 25% of the observations be below?
1
Expert's answer
2019-12-09T11:52:54-0500

a)


"P(z>z')=0.15\to P(z<z')=0.85"

From z-table:


"z'=1.04"

"x'=20+(1.04)4=24.16"

b)


"P(z>z')=0.75\to P(z<z')=0.25"

From z-table:


"z'=-0.67"

"x'=20+(-0.67)4=17.32"

c)


"P(z<-z')=0.2, P(z>z')=0.2"

From z-table:


"P(z<-0.84)=0.2005"

"z'=0.84"

"x'_1=20+(-0.84)4=16.64"

"x'_2=20+(0.84)4=23.36"

d)


"P(z<z')=0.25"

Thus, this is the same value as in the part b).


"x'=17.32"